OpenStudy (anonymous):

how do you explain to solve a2 + 12ab + 27b2

OpenStudy (whpalmer4):

\[a^2+12ab+27b^2\] (It really isn't difficult to use the equation entry form, and it makes the problems much clearer for those trying to help you!) I assume you are trying to factor this expression. To factor, notice that the final term has a coefficient of 27, which can be created by 3*9, and the middle term a coefficient of 12, which can be created by 3+9. If our equation can be written as \[x^2+(n+m)xy+(n*m)y^2\] then the factoring is \[x^2+(n+m)xy+(n*m)y^2 = (x+n)(x+m)\] We're in luck! We can write our equation as \[a^2+12ab+27b^2 = (a+3b)(a+9b)\] Set (a+3b)=0 or (a+9b)=0 to find the solutions.